the scale of a map is 1/2 in.:8 mi. i need to find the actual distance for each map distance 2 in. if you just teach me how to do this one i can do the rest on my own !
Answer by user_dude2008(1862)

(1/2)/(8) = 2/x


(1/2)x = 2*8


(1/2)x = 16


x = 2*16

x = 32

Answer: 2 in. corresponds to 32 mi.